OpenStudy (anonymous):

The supplement of an angle is 3 times the complement of an angle. Find half of the complement of the angle.

OpenStudy (ajprincess):

Let the angle be x and the compliment of that angle be y. Sum of complimentary angles=90 x+y=90 x=90-y Let the supplement of x be z. z=3y Sum of supplementary angles=180 x+z=180 x+3y=180 90-y+3y=180 Can u find y?

OpenStudy (ajprincess):

@anab